# ACS CONDITION ON MINIMAL ISOPARAMETRIC HYPERSURFACES OF POSITIVE RICCI CURVATURE IN UNIT SPHERES

### 10 Mar 2026 v1

@@ Abstract.
Motivated by the Schoen–Marques–Neves conjecture relating the Morse index of a minimal hypersurface to its first Betti number in ambient manifolds with positive Ricci curvature, we study the Ambrozio–Carlotto–Sharp (ACS) criterion for index growth. We verify a sufficient pointwise ACS inequality for several families of minimal isoparametric hypersurfaces with positive Ricci curvature in the unit sphere. As a consequence, for any closed embedded minimal hypersurface $M^{n}$ in such an ambient manifold $\mathcal{N}^{n+1}\subset\mathbb{S}^{n+2}\subset\mathbb{R}^{n+3}$, one has $\operatorname{index}(M)\geq\frac{2}{d(d-1)}\,b_{1}(M),d=n+3,$ where $b_{1}(M)$ is the first Betti number of $M$.
@@ 1. Introduction
    If $g=3$ and $m_{1}=m_{2}=4$ or $8$, then $N$ satisfies the ACS condition.
@@
    If $g=4$ and $m_{1},m_{2}\geq 5$, then $N$ satisfies the ACS condition.
@@
Gorodski, Mendes and Radeschi [6] also produced many examples supporting the Schoen-Marques-Neves conjecture. They obtained robust index lower bounds for minimal hypersurfaces in several isoparametric and symmetric ambient spaces by a different method. Our contribution is instead to verify a sufficient ACS inequality for the ambient manifold itself.Our arguments in the $g=4$ case are deliberately coarse: we impose a uniform bound on the largest absolute principal curvature. This yields a clean sufficient condition $\min\{m_{1},m_{2}\}\geq 5$.
@@
    for $g=4$, our estimate does not decide the ACS condition when at least one multiplicity lies in $\{2,3,4\}$;
@@ 3. The computations
We next compute the second fundamental form terms. Since
$\{e_{1},\dots,e_{n},\nu\}$ is an orthonormal basis of $TN$,
@@
$\Delta:={}$ $(2n-2)-2\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}^{2}-2\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}^{2}+2\Big(\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i}\Big)^{2}$
By Proposition 3.1, it suffices to show that $\Delta>0$.
@@
If the common multiplicity is $m=4$ or $8$, then
@@
Therefore $N$ satisfies the ACS condition when $g=3$ and $m=4$ or $8$.
@@
$\sum_{i}x_{i}^{2}\lambda_{i}^{2}\leq a^{2},\qquad\sum_{i}\nu_{i}^{2}\lambda_{i}^{2}\leq a^{2},\qquad\left|\sum_{i}x_{i}^{2}\lambda_{i}\right|\leq a,\qquad\left|\sum_{i}\nu_{i}^{2}\lambda_{i}\right|\leq a,$
Proposition 3.1 gives
@@
Condition (9) is therefore equivalent to
@@
Equivalently, if $m_{1},m_{2}\geq 5$, then (9) holds, and therefore (8) yields $\Delta>0$. This proves the ACS condition for the case $g=4$ with $m_{1},m_{2}\geq 5$.
The index estimate then follows immediately from Theorem 1.1. ∎
@@
The argument above leaves the following cases unresolved by this method: the case $g=3$ with multiplicity $m=2$, and the case $g=4$ when at least one multiplicity belongs to $\{2,3,4\}$. The cases $g=3$ with $m=1$ and $g=6$ are excluded because the ambient Ricci curvature is not positive.

### 23 Mar 2026 v2

@@ Abstract.
We study the Ambrozio–Carlotto–Sharp (ACS) criterion on minimal isoparametric hypersurfaces $N^{n+1}\subset S^{n+2}$ with positive Ricci curvature, motivated by the Schoen–Marques–Neves conjecture on Morse index. For $g=4$ distinct principal curvatures with multiplicities $m_{1},m_{2}$, we prove that the pointwise ACS inequality holds if and only if $\min\{m_{1},m_{2}\}\geq 4$. Sufficiency is obtained via a moment-relaxation technique yielding the sharp bound $4a^{2}$ on the quadratic part of the integrand; necessity follows from an explicit extremal configuration in the top eigenspace of the shape operator. We also verify the ACS condition for $g=3$ with $m_{1}=m_{2}\in\{4,8\}$. As a consequence, for any closed embedded minimal hypersurface $M^{n}$ in such an ambient manifold, $\operatorname{index}(M)\geq\tfrac{2}{d(d-1)}\,b_{1}(M)$ with $d=n+3$.
@@ 1. Introduction
    If $g=3$ and $m_{1}=m_{2}\in\{4,8\}$, then $N$ satisfies the ACS condition.
@@
    If $g=4$ and $\min\{m_{1},m_{2}\}\geq 4,$ then $N$ satisfies the ACS condition.
@@
Gorodski, Mendes and Radeschi [6] also produced many examples supporting the Schoen-Marques-Neves conjecture. They obtained robust index lower bounds for minimal hypersurfaces in several isoparametric and symmetric ambient spaces by a different method. Our contribution is instead to verify a sufficient ACS inequality for the ambient manifold itself. In the $g=4$ case, we work with the pointwise quantity $\Delta(X,\nu)$ and obtain progressively tighter bounds. A direct estimate of $\Delta$ using only the largest absolute curvature yields the range $\min\{m_{1},m_{2}\}\geq 5$. By retaining a quadratic term in Proposition 3.1 and completing the square, we lower the coefficient to $\tfrac{17}{4}$, which extends the result to $\min\{m_{1},m_{2}\}=4$ with $m_{1}+m_{2}\leq 19$. Finally, a moment-relaxation technique (Proposition 3.2) that exploits the asymmetry between the largest and smallest principal curvatures produces the sharp coefficient $4$, establishing the ACS condition for all pairs with $\min\{m_{1},m_{2}\}\geq 4$ without restriction on $m_{1}+m_{2}$. We observe in Remark 3.4 that $\min\{m_{1},m_{2}\}\geq 4$ is also a necessary condition for any argument based solely on pointwise positivity of $\Delta$. Thus our result is sharp with respect to this method.
@@
    for $g=4$, our estimate does not decide the ACS condition when at least one multiplicity lies in $\{2,3\}$;
@@ 3. The computations
We next compute the second fundamental form terms. Since $\{e_{1},\dots,e_{n},\nu\}$ is an orthonormal basis of $TN$,
@@
Proposition 3.2.
Assume $g=4$, set
$q:=\min\{m_{1},m_{2}\}=4,\qquad s:=m_{1}+m_{2},$
and let
$a^{2}=\frac{1+\sqrt{1-4/s}}{1-\sqrt{1-4/s}}.$
If $s\geq 15$, then for every orthogonal unit vectors $X,\nu\in T_{p}N$ one has
$\Delta(X,\nu)\geq 4s-4-4a^{2}>0.$
In particular, $N$ satisfies the ACS condition.
Proof.
After interchanging $m_{1}$ and $m_{2}$ if necessary, we may assume
$m_{1}=4\leq m_{2}.$
For the minimal leaf with $g=4$, we have
$\cos 2\theta_{1}=\sqrt{\frac{m_{2}}{s}},\qquad\sin 2\theta_{1}=\sqrt{\frac{m_{1}}{s}}=\sqrt{\frac{4}{s}},$
and therefore
$\lambda_{1}^{2}=\frac{1+\sqrt{m_{2}/s}}{1-\sqrt{m_{2}/s}},\qquad\lambda_{4}^{2}=\frac{1+\sqrt{m_{1}/s}}{1-\sqrt{m_{1}/s}}.$
Since $m_{2}\geq m_{1}$ and the function
$t\longmapsto\frac{1+\sqrt{t}}{1-\sqrt{t}}$
is strictly increasing on $(0,1)$, it follows that $\lambda_{1}^{2}\geq\lambda_{4}^{2}$. By definition of $a$, we therefore have
$a=\lambda_{1}=\cot\theta_{1}>0.$
Moreover, $\lambda_{1}$ has multiplicity $m_{1}=4$, so the $a$-eigenspace has dimension exactly $4$.
Now write $\lambda_{4}=-b<0$. Then
$b:=-\lambda_{4}=-\cot\Bigl(\theta_{1}+\frac{3\pi}{4}\Bigr)=\tan\Bigl(\theta_{1}+\frac{\pi}{4}\Bigr)=\frac{a+1}{a-1}.$
Hence all principal curvatures lie in the interval [b,a][-b,a].
Next we explain why $b\leq a/2$ when $s\geq 15$. Set
$x(s):=\sqrt{1-\frac{4}{s}}\in(0,1).$
Then
$a^{2}=\frac{1+x(s)}{1-x(s)}.$
Since $x(s)$ is increasing in $s$ and the function $x\mapsto\frac{1+x}{1-x}$ is strictly increasing on $(0,1)$, the quantity $a^{2}$ is increasing in $s$. Hence the assumption $s\geq 15$ gives
$a^{2}\geq\frac{1+\sqrt{11/15}}{1-\sqrt{11/15}}=\frac{13+\sqrt{165}}{2}>\frac{13+3\sqrt{17}}{2}=\left(\frac{3+\sqrt{17}}{2}\right)^{2}.$
Therefore
$a>\frac{3+\sqrt{17}}{2}.$
On the other hand, since $a>1$ and $b=\frac{a+1}{a-1}$, we have
$b\leq\frac{a}{2}\iff\frac{a+1}{a-1}\leq\frac{a}{2}\iff 2(a+1)\leq a(a-1)\iff a^{2}-3a-2\geq 0\iff a\geq\frac{3+\sqrt{17}}{2}.$
Thus the previous inequality for $a$ implies
$b\leq\frac{a}{2}.$
Now define
$A:=\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i},\qquad B:=\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i},\qquad C:=\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i},$
and
$S_{x}:=\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}^{2},\qquad S_{\nu}:=\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}^{2}.$
By Proposition 3.1,
$\Delta=(2n-2)-2S_{x}-2S_{\nu}+2C^{2}+B^{2}-AB.$
Equivalently,
$\Delta=(2n-2)-F+2C^{2},\qquad F:=2S_{x}+2S_{\nu}-B^{2}+AB.$
Since each principal curvature $\lambda_{i}$ lies in [b,a][-b,a], we have
$(\lambda_{i}-a)(\lambda_{i}+b)\leq 0,$
hence
$\lambda_{i}^{2}\leq(a-b)\lambda_{i}+ab.$
Averaging against the weights $x_{i}^{2}$ and $\nu_{i}^{2}$, we obtain
$S_{x}\leq(a-b)A+ab,\qquad S_{\nu}\leq(a-b)B+ab.$
Therefore
$F\leq G(A,B),$
where
$G(A,B):=2(a-b)A+2(a-b)B+4ab-B^{2}+AB.$
Since $A,B\in[-b,a]$, we have
$\frac{\partial G}{\partial A}=2(a-b)+B\geq 2(a-b)-b=2a-3b\geq\frac{a}{2}>0,$
because $b\leq a/2$. Thus $G$ is strictly increasing in $A$ throughout the square $b,a]\times[-b,a]$, and so its maximum is attained at $A=a$. Hence
$G(A,B)\leq G(a,B)=-B^{2}+(3a-2b)B+2a^{2}+2ab.$
This is a concave quadratic in $B$, with axis
$B_{0}=\frac{3a-2b}{2}=\frac{3a}{2}-b\geq a,$
since $b\leq a/2$. It follows that $G(a,B)$ is increasing on [b,a][-b,a], and therefore
$F\leq G(A,B)\leq G(a,a)=4a^{2}.$
Consequently,
$\Delta(X,\nu)\geq(2n-2)-4a^{2}.$
Since n=2s1n=2s-1, this becomes
$\Delta(X,\nu)\geq 4s-4-4a^{2}.$
Finally, let
$x:=\sqrt{1-\frac{4}{s}}\in(0,1).$
Then
$a^{2}=\frac{1+x}{1-x},\qquad s=\frac{4}{1-x^{2}},$
so
$s-1-a^{2}=\frac{3+x^{2}}{1-x^{2}}-\frac{1+x}{1-x}=\frac{2-2x}{1-x^{2}}=\frac{2}{1+x}>0.$
Hence
$4s-4-4a^{2}=4(s-1-a^{2})>0,$
and therefore $\Delta(X,\nu)>0$ for every orthogonal unit pair $(X,\nu)$.
Moreover, this lower bound is sharp. Indeed, the $a$-eigenspace has dimension $q=4$, so we may choose orthogonal unit vectors $X,\nu$ in that eigenspace. Then
$A=B=a,\qquad S_{x}=S_{\nu}=a^{2},\qquad C=0,$
and Proposition 3.1 gives
$\Delta(X,\nu)=4s-4-4a^{2}.$
@@
$A:=\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i},\qquad B:=\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i},\qquad C:=\sum_{i=1}^{n+1}x_{i}\nu_{i}\lambda_{i},$
and
$S_{x}:=\sum_{i=1}^{n+1}x_{i}^{2}\lambda_{i}^{2},\qquad S_{\nu}:=\sum_{i=1}^{n+1}\nu_{i}^{2}\lambda_{i}^{2}.$
Then Proposition 3.1 gives
$\Delta=(2n-2)-2S_{x}-2S_{\nu}+2C^{2}+B^{2}-AB.$
It suffices to show that $\Delta>0$.
@@
If the common multiplicity is $m\in\{4,8\}$, then
@@
Therefore $N$ satisfies the ACS condition when $g=3$ and $m\in\{4,8\}$.
@@
$S_{x}\leq a^{2},\qquad S_{\nu}\leq a^{2},\qquad|A|\leq a,\qquad|B|\leq a,$
we first record the rough estimate obtained by discarding the nonnegative terms $2C^{2}$ and $B^{2}$ and bounding $|AB|\leq a^{2}$:
@@
The rough estimate (8) gives the range $q\geq 5$ as follows. Condition (9) is equivalent to
@@
To handle the borderline case $q=4$, we keep the term $B^{2}$ instead of discarding it. Since $2C^{2}\geq 0$,
$\Delta\geq(2n-2)-2S_{x}-2S_{\nu}+B^{2}-AB.$
Completing the square yields
$B^{2}-AB=\left(B-\frac{A}{2}\right)^{2}-\frac{A^{2}}{4}\geq-\frac{A^{2}}{4}.$
Moreover, since $\sum_{i}x_{i}^{2}=1$, the Cauchy–Schwarz inequality gives
$A^{2}=\left(\sum_{i}x_{i}^{2}\lambda_{i}\right)^{2}\leq\sum_{i}x_{i}^{2}\lambda_{i}^{2}=S_{x}.$
Therefore
(11) $\Delta\geq(2n-2)-\tfrac{9}{4}\,S_{x}-2S_{\nu}\geq(2n-2)-\tfrac{17}{4}\,a^{2}.$
Hence it is enough to require
(12) $a^{2}<\frac{4}{17}(2n-2)=\frac{16}{17}(s-1).$
If $q=4$, then p=s4p=s-4, so
$a^{2}=\frac{1+\sqrt{1-4/s}}{1-\sqrt{1-4/s}}.$
Substituting this into (12) and simplifying, we obtain
$16s^{2}-304s-1<0.$
For integers $s$, this holds exactly when $s\leq 19$: the left-hand side is increasing for $s\geq 10$, equals 1-1 at $s=19$, and equals $319$ at $s=20$. Therefore (11) yields $\Delta>0$ whenever $q=4$ and $s\leq 19$. On the other hand, for $q=4$ and $s\geq 15$, Proposition 3.2 gives
$\Delta(X,\nu)\geq 4s-4-4a^{2}>0.$
Since the ranges $s\leq 19$ and $s\geq 15$ cover all $s\geq 8$, the ACS condition holds for every multiplicity pair with $q=4$.
This proves the ACS condition in all cases listed in Theorem 1.2. The index estimate then follows immediately from Theorem $1.$ ∎
@@
The argument above leaves the following cases unresolved by this method: the case $g=3$ with multiplicity $m=2$, and the case $g=4$ when at least one multiplicity belongs to $\{2,3\}$. The cases $g=3$ with $m=1$ and $g=6$ are excluded because the ambient Ricci curvature is not positive.
Remark 3.4.
The bound $\min\{m_{1},m_{2}\}\geq 4$ in Theorem 1.2(2) is sharp with respect to the pointwise method. Indeed, the principal curvature whose absolute value equals $a$ has multiplicity $q:=\min\{m_{1},m_{2}\}$. Since positive Ricci curvature forces $q\geq 2$ in the case $g=4$, the corresponding eigenspace has dimension $\geq 2$, and we may choose orthogonal unit vectors $X,\nu$ supported entirely in that eigenspace. For such a choice,
$C=0,\qquad S_{x}=S_{\nu}=a^{2},\qquad A=B=\lambda\quad(\text{where }|\lambda|=a),$
and the terms $B^{2}-AB$ and $2C^{2}$ all vanish. Proposition 3.1 therefore gives
$\Delta(X,\nu)=(2n-2)-4a^{2}.$
Since pointwise positivity requires $\Delta(X,\nu)>0$ for every orthogonal unit pair $(X,\nu)$, and the configuration above attains the value $(2n-2)-4a^{2}$, such positivity forces $(2n-2)-4a^{2}>0$. Substituting
$a^{2}=\frac{1+\sqrt{1-q/s}}{1-\sqrt{1-q/s}},\qquad s=m_{1}+m_{2},$
and simplifying, this condition is equivalent to q>44/sq>4-4/s, and hence to $q\geq 4$ for integer $q$ and $s\geq 4$. Thus $\min\{m_{1},m_{2}\}\geq 4$ is a necessary condition for any argument based only on pointwise positivity of $\Delta$.
In the present paper, this necessary condition is also shown to be sufficient. Indeed, the completed-square estimate (11) covers the range $8\leq s\leq 19$, while Proposition 3.2 proves that for $s\geq 15$ the moment-relaxation bound $F\leq 4a^{2}$ is achieved precisely at the extremal configuration above. Since
$4(s-1)-4a^{2}=\frac{8}{1+\sqrt{1-4/s}}>0,$
the argument closes for every $s\geq 8$, i.e. for every multiplicity pair with $q=4$.